OpenStudy (anonymous):

Michael has an average of 89 on the five test he has taken so far this semester.what is the minimum grade michael must earn on the next test to earn an average grade of 90?

OpenStudy (anonymous):

@hyntc

OpenStudy (anonymous):

Alright so he got a total of 356 points in the last tests(average=total 4 tests points/4) I just multiplied by 4. (356+x)/5=90 450=356+x x=94

OpenStudy (anonymous):

445 +x/6=90 <----- That would be the average of the sum of the 5 tests plus X divided by 6 tests. Then you would need to isolate x by multiplying both sides by 6 445 + x=540 Then subtract both sides by 445 x = 540-445 x = 95